### v2

@@ 2.4.2 Adversarial robustness
Strategic borrowers and lenders seek to manipulate the protocol’s interest rate to their advantage by exploiting the structure of the interest rate controller. They may simulate the protocol’s responses to various demand and supply scenarios, selecting the one that maximizes their own utility rather than acting truthfully. We aim to quantify how much influence these strategic users can exert on the interest rate. Specifically, we consider a lending pool governed by the protocol $\mathcal{P}$ with a set of truthful users characterized by fixed demand and supply curves $f(r;\theta)$ and $g(r\,U;\omega)$; Plus one strategic lender, $\mathcal{A}^{l}$, who has an external interest rate $r^{l}$ and a deposit of $\delta_{l}\times\big{(}\max_{x}g(x;\omega)\big{)}$, and one strategic borrower, $\mathcal{A}^{b}$, with an external rate $r^{b}$ and demand of $\delta_{b}\times\big{(}\max_{x}f(x;\theta)\big{)}$.
@@
$AI^{\mathcal{P}}(\delta_{b},\delta_{l},\theta,\omega)\coloneqq\sup_{r^{b},r^{l}}\mathbb{E}\left[\left|r_{\text{truthful}}-r_{\text{strategic}}\right|\right]$ (7)
For ease of notation we also denote adversarial impact by $AI^{\mathcal{P}}(\delta_{b},\delta_{l})$.
2.4.3 Risk Control
@@
Pool defaults occur when a borrower’s debt exceeds the value of their collateral due to a price drop, making the pool unable to recover the full loan amount. Liquidation occurs when a borrower, maintaining the maximum loan-to-value ratio, faces a price drop that forces the protocol to liquidate part of their collateral to bring the ratio back below the liquidation threshold. These risks are formally defined in Appendix [0.A.7], where the exact conditions for pool default and liquidation are detailed.
@@ 2.5 Baseline
$r_{t}=\begin{cases}R_{\text{slope1}}\,\frac{U_{t}}{U^{*}},&\text{if }U_{t}\leq U^{*}\\ R_{\text{slope1}}+R_{\text{slope2}}\left(\frac{U_{t}-U^{*}}{1-U^{*}}\right),&\text{if }U_{t}>U^{*}\end{cases}$ (8)
@@
3 Protocol design
@@
$B$ =$=\begin{cases}L_{t},&\qquad\quad\text{if }r_{t}<r_{\text{min}},\\ -\mathtt{a}_{t}^{b}r_{t}+\mathtt{b}_{t}^{b},&\qquad\quad\text{if }r_{\text{min}}\leq r_{t}\leq r_{\text{max}},\\ 0,&\qquad\quad\text{if }r_{t}>r_{\text{max}},\end{cases}$ (9)
@@
$L$ =$=\begin{cases}B_{t},&\qquad\quad\text{if }r_{t}U_{t}<r_{\text{min}},\\ \mathtt{a}_{t}^{l}r_{t}U_{t}-\mathtt{b}_{t}^{l},&\qquad\quad\text{if }r_{\text{min}}\leq r_{t}U_{t}\leq r_{\text{max}},\\ \infty,&\qquad\quad\text{if }r_{t}U_{t}>r_{\text{max}},\end{cases}$ (10)
@@
Estimating the parameters Our proposed interest rate controller employs a Recursive Least Squares (RLS) estimator with a forgetting factor to adaptively estimate the parameters $\mathtt{a}_{t}^{b}$, $\mathtt{b}_{t}^{b}$, $\mathtt{a}_{t}^{l}$, and $\mathtt{b}_{t}^{l}$. The RLS with a forgetting factor is a recursive version of the least squares estimator designed to handle systems modeled by $y_{t}=\mathbf{x}_{t}^{T}\bm{\theta}_{t}+\varepsilon_{t}$, where $\bm{\theta}_{t}$ is the state of the system that can change arbitrarily over time; And $\varepsilon_{t}$ is an independent, zero-mean noise term with variance $\nu$. Given a new sample $(\mathbf{x}_{t},y_{t})$ at any time $t$, the RLS estimator updates its estimate of $\bm{\theta}_{t}$, denoted as $\hat{\bm{\theta}}_{t}$, by minimizing the following cost function [26, 14]:
$J_{t}(\bm{\theta})=\sum_{\tau=0}^{t}\rho^{t-\tau}(y_{\tau}-\mathbf{x}_{\tau}^{T}\bm{\theta})^{2},$ (11)
@@
1:Initialize: $\hat{\bm{\theta}}^{l}_{0},\hat{\bm{\theta}}^{b}_{0}$, $\mathbf{P}^{l}_{0},\mathbf{P}^{b}_{0}\leftarrow$ large positive definite matrix, $\rho$ (forgetting factor), $\nu$ noise standard deviation
2:for each time step $t$ do
3:     Observe $r_{t-1},U_{t-1}$ and $L_{t},B_{t}$; And set $\mathbf{x}_{t}^{b}=[r_{t-1},1]^{T}$ , $\mathbf{x}_{t}^{l}=[r_{t-1}U_{t-1},1]^{T}$
4:     Compute the gain vector for the supply and demand observation :
$\mathbf{K}_{t}^{l}=\frac{\mathbf{P}^{l}_{t-1}\mathbf{x}^{l}_{t}}{(\rho\,\nu^{2}+(\mathbf{x}^{l}_{t})^{T}\mathbf{P}^{l}_{t-1}\mathbf{x}^{l}_{t})},\quad\mathbf{K}_{t}^{b}=\frac{\mathbf{P}^{b}_{t-1}\mathbf{x}^{b}_{t}}{(\rho\,\nu^{2}+(\mathbf{x}^{b}_{t})^{T}\mathbf{P}^{b}_{t-1}\mathbf{x}^{b}_{t})}$
5:     Update the parameter estimate:
@@
6:     Update the covariance matrix:
@@
7:     Parse $[\hat{\mathtt{a}}_{t}^{l},-\hat{\mathtt{b}}_{t}^{l}]\leftarrow\hat{\bm{\theta}}_{t}^{l}$ and $[-\hat{\mathtt{a}}_{t}^{b},\hat{\mathtt{b}}_{t}^{b}]\leftarrow\hat{\bm{\theta}}_{t}^{b}$
8:end for
@@
Optimizing interest rate for desired utilization Given the demand and supply models presented in Equations 9 and 10, when the noise is small, the utilization rate can be approximated with:
$U_{t+1}\approx\begin{cases}1&\text{if }r_{t}<r_{\text{min}},\\ \frac{-\mathtt{a}_{t}^{b}r_{t}+\mathtt{b}_{t}^{b}}{\mathtt{a}_{t}^{l}r_{t}U_{t}-\mathtt{b}_{t}^{l}}&\text{if }r_{\text{min}}\leq r_{t},\text{ and }r_{t}U_{t}\geq r_{\text{max}},\\ 0&\text{if }r_{t}U_{t}>r_{\text{max}},\\ \end{cases}$ (12)
The objective of our protocol is to set the interest rate $r_{t}$ such that the utilization remains close to a desired utilization $U^{*}$. The optimal interest rate $r^{*}_{\text{utl}}$ that achieves this desired utilization can be derived as follows:  [^4] We develop similar results for the objective of maximizing revenue instead of optimizing rate deviation. See Appendix [0.A.4]
$\frac{-\mathtt{a}_{t}^{b}\,r^{*}_{\text{utl}}+\mathtt{b}_{t}^{b}}{\mathtt{a}_{t}^{l}r^{*}_{\text{utl}}\,U^{*}-\mathtt{b}_{t}^{l}}=U^{*}\implies r^{*}_{\text{utl}}=\frac{\mathtt{b}_{t}^{b}+\mathtt{b}_{t}^{l}U^{*}}{\mathtt{a}_{t}^{b}+\mathtt{a}_{t}^{l}(U^{*})^{2}}$ (13)
@@
1:Initialize: $U^{*}$, $\zeta$ (randomness probability), $\xi$ (error threshold to do randomization)
2:for each time step $t$ do
3:     Read $\hat{\mathtt{a}}_{t}^{b},\hat{\mathtt{b}}_{t}^{b},\hat{\mathtt{a}}_{t}^{l},\hat{\mathtt{b}}_{t}^{l},\mathbf{P}^{l}_{t},\mathbf{P}^{b}_{t}$ from Algorithm 1
4:     $\text{Var}(\hat{\mathtt{a}}_{t}^{b})\leftarrow\mathbf{P}^{b}_{t}(1,1)$, $\text{Var}(\hat{\mathtt{b}}_{t}^{b})\leftarrow\mathbf{P}^{b}_{t}(2,2)$, $\text{Var}(\hat{\mathtt{a}}_{t}^{l})\leftarrow\mathbf{P}^{l}_{t}(1,1)$, $\text{Var}(\hat{\mathtt{b}}_{t}^{l})\leftarrow\mathbf{P}^{l}_{t}(2,2)$
5:     $\mathbb{E}\left[r_{t}\right]\leftarrow\frac{\hat{\mathtt{b}}_{t}^{b}+\hat{\mathtt{b}}_{t}^{l}U^{*}}{\hat{\mathtt{a}}_{t}^{b}+\hat{\mathtt{a}}_{t}^{l}(U^{*})^{2}}$
6:     $\text{Var}(r_{t})\leftarrow\frac{1}{\big{(}\hat{\mathtt{a}}_{t}^{b}+\hat{\mathtt{a}}_{t}^{l}(U^{*})^{2}\big{)}^{2}}\left(\text{Var}(\hat{\mathtt{b}}_{t}^{b})+(U^{*})^{2}\text{Var}(\hat{\mathtt{b}}_{t}^{l})\right)$
7:$\quad+\frac{\big{(}\hat{\mathtt{b}}_{t}^{b}+\hat{\mathtt{b}}_{t}^{l}U^{*}\big{)}^{2}}{\big{(}\hat{\mathtt{a}}_{t}^{b}+\hat{\mathtt{a}}_{t}^{l}(U^{*})^{2}\big{)}^{4}}\left(\text{Var}(\hat{\mathtt{a}}_{t}^{b})+(U^{*})^{4}\text{Var}(\hat{\mathtt{a}}_{t}^{l})\right)$
8:     Sample $r_{t}\sim\mathbb{N}(\mathbb{E}\left[r_{t}\right],\text{Var}[r_{t}])$
9:end for
@@
with probability at least $1-\delta$, where $\limsup_{t\to\infty}\psi_{t}(\rho)=\frac{\nu^{2}}{\rho^{2N}\ln\left(\frac{1}{\rho}\right)}$ and $N$ is given by: $N=\Theta\left(\frac{\ln 1/\delta}{\ln\min_{t}\left(\text{diam}(\mathbf{P}^{b}_{t}+\mathbf{P}^{l}_{t})\right)}\right).$
@@
Intuitively, the lemma presented above provides a method for adjusting the liquidation threshold $LT_{t}$ and the collateral factor $c_{t}$ to minimize these risks. By setting $LT_{t}$ appropriately, the protocol can keep expected defaults below a threshold, while also ensuring that liquidations remain below a certain threshold. The relationship between the loan-to-value ratio and the liquidation threshold is key for dynamically adjusting risk parameters. This approach is based on [13], where $LT_{t}$ serves as an upper bound for the loan-to-value ratio, helping to limit the likelihood of default.
4 Adversarial analysis
@@
4.1 Learning-based interest rate controller
Learning-based interest rate controllers estimate demand and supply while incorporating prior beliefs to set optimal rates. As adaptivity increases, these controllers become more vulnerable to manipulation, quickly updating beliefs based on a few manipulated samples. To quantify this impact, we analyze an adaptive learning-based algorithm, $\mathtt{LC}$, which alternates between exploration and exploitation phases. During exploration, the controller detects discrepancies between observed and expected demand and supply, testing various rates to update its beliefs. Once stabilized, the controller enters exploitation, setting rates based on the learned functions. This cycle repeats as new changes are detected. We model $\mathtt{LC}$ as an auction-like mechanism. In exploration, borrowers and lenders report their minimum (or maximum) acceptable rates and quantities. $\mathtt{LC}$ infers the demand and supply curves from this data. In exploitation, it sets a single rate $r_{t}$, allowing lenders with bids $r_{t}$ or lower to deposit, and borrowers with bids $r_{t}$ or higher to receive loans. This mimics a uniform price auction, where participants’ reported bids influence the interest rate, maximizing utility based on private valuations. Participant utility depends on whether the rate falls within their acceptable range and the portion of their supply or demand allocated to the protocol or external markets. More details on this auction implementation are provided in Appendix [0.A.5].

### v3

@@ 3.4.2. Adversarial robustness
Strategic borrowers and lenders seek to manipulate the protocol’s interest rate to their advantage by exploiting the structure of the interest rate controller. They may simulate the protocol’s responses to various demand and supply scenarios, selecting the one that maximizes their own utility rather than acting truthfully. We aim to quantify how much influence these strategic users can exert on the interest rate. Specifically, we consider a lending pool governed by the protocol $\mathcal{P}$ with a set of truthful users characterized by fixed demand and supply curves $f(r;\mathcal{B}_{t})$ and $g(r_{\text{eff}};\mathcal{L}_{t})$; Plus one strategic lender, $\mathcal{A}^{l}$, who has an external interest rate $r^{l}$ and a deposit of $\delta_{l}\times\big{(}\max_{x}g(x;\mathcal{L}_{t})\big{)}$, and one strategic borrower, $\mathcal{A}^{b}$, with an external rate $r^{b}$ and demand of $\delta_{b}\times\big{(}\max_{x}f(x;\mathcal{B}_{t})\big{)}$.
@@
(9) $AI^{\mathcal{P}}(\delta_{b},\delta_{l})\coloneqq\sup_{r^{b},r^{l}}\mathbb{E}\left[\left|r_{\text{truthful}}-r_{\text{strategic}}\right|\right]$
3.4.3. Liquidation and default risk control
@@
Pool defaults occur when a borrower’s debt exceeds the value of their collateral due to a price drop, making the pool unable to recover the full loan amount. Liquidation occurs when a borrower, maintaining the maximum loan-to-value ratio, faces a price drop that forces the protocol to liquidate part of their collateral to bring the ratio back below the liquidation threshold. These risks are formally defined in Appendix [A.5], where the exact conditions for pool default and liquidation are detailed.
@@ 3.5. Baseline
(10) $r_{t}=\begin{cases}R_{\text{slope1}}\,\frac{U_{t}}{U^{*}},&\text{if }U_{t}\leq U^{*}\\ R_{\text{slope1}}+R_{\text{slope2}}\left(\frac{U_{t}-U^{*}}{1-U^{*}}\right),&\text{if }U_{t}>U^{*}\end{cases}$
@@
4. Protocol design
4.1. Interest rate controller
@@
(11) $B$ =$=\begin{cases}L_{t},&\qquad\quad\text{if }r_{t}<r_{\text{min}},\\ -\mathtt{a}_{t}^{b}r_{t}+\mathtt{b}_{t}^{b},&\qquad\quad\text{if }r_{\text{min}}\leq r_{t}\leq r_{\text{max}},\\ 0,&\qquad\quad\text{if }r_{t}>r_{\text{max}},\end{cases}$
@@
(12) $L$ =$=\begin{cases}B_{t},&\qquad\quad\text{if }r_{t}U_{t}<r_{\text{min}},\\ \mathtt{a}_{t}^{l}r_{t}U_{t}-\mathtt{b}_{t}^{l},&\qquad\quad\text{if }r_{\text{min}}\leq r_{t}U_{t}\leq r_{\text{max}},\\ \infty,&\qquad\quad\text{if }r_{t}U_{t}>r_{\text{max}},\end{cases}$
@@
Estimating the parameters Our proposed interest rate controller employs a Recursive Least Squares (RLS) estimator with a forgetting factor to adaptively estimate the parameters $\mathtt{a}_{t}^{b}$, $\mathtt{b}_{t}^{b}$, $\mathtt{a}_{t}^{l}$, and $\mathtt{b}_{t}^{l}$. The RLS with a forgetting factor is a recursive version of the least squares estimator designed to handle systems modeled by $y_{t}=\mathbf{x}_{t}^{T}\bm{\theta}_{t}+\varepsilon_{t}$, where $\bm{\theta}_{t}$ is the state of the system that can change arbitrarily over time; And $\varepsilon_{t}$ is an independent, zero-mean noise term. Given a new sample $(\mathbf{x}_{t},y_{t})$ at any time $t$, the RLS estimator updates its estimate of $\bm{\theta}_{t}$, denoted as $\hat{\bm{\theta}}_{t}$, by minimizing the following cost function (Islam and Bernstein, 2019; Canetti and España, 1989):
(13) $J_{t}(\bm{\theta})=\sum_{\tau=0}^{t}\rho^{t-\tau}(y_{\tau}-\mathbf{x}_{\tau}^{T}\bm{\theta})^{2},$
@@
Initialize: $\hat{\bm{\theta}}^{l}_{0},\hat{\bm{\theta}}^{b}_{0}$, $\mathbf{P}^{l}_{0},\mathbf{P}^{b}_{0}\leftarrow$ large positive definite matrix, $\rho$ (forgetting factor), $\nu$ noise standard deviation for each time step $t$ do
Observe $r_{t-1},U_{t-1}$ and $L_{t},B_{t}$; And set $\mathbf{x}_{t}^{b}=[r_{t-1},1]^{T}$ , $\mathbf{x}_{t}^{l}=[r_{t-1}U_{t-1},1]^{T}$
Compute the gain vector for the supply and demand observation :
$\mathbf{K}_{t}^{l}=\frac{\mathbf{P}^{l}_{t-1}\mathbf{x}^{l}_{t}}{(\rho+(\mathbf{x}^{l}_{t})^{T}\mathbf{P}^{l}_{t-1}\mathbf{x}^{l}_{t})},\quad\mathbf{K}_{t}^{b}=\frac{\mathbf{P}^{b}_{t-1}\mathbf{x}^{b}_{t}}{(\rho+(\mathbf{x}^{b}_{t})^{T}\mathbf{P}^{b}_{t-1}\mathbf{x}^{b}_{t})}$
Update the parameter estimate:
@@
Update the covariance matrix:
@@
      Parse $[\hat{\mathtt{a}}_{t}^{l},-\hat{\mathtt{b}}_{t}^{l}]\leftarrow\hat{\bm{\theta}}_{t}^{l}$ and $[-\hat{\mathtt{a}}_{t}^{b},\hat{\mathtt{b}}_{t}^{b}]\leftarrow\hat{\bm{\theta}}_{t}^{b}$
end for
@@
Optimizing interest rate for desired utilization Given the demand and supply models presented in Equations 11 and 12, when the noise is small, the utilization rate can be approximated with:
(14) $U_{t+1}\approx\begin{cases}1&\text{if }r_{t}<r_{\text{min}},\\ \frac{-\mathtt{a}_{t}^{b}r_{t}+\mathtt{b}_{t}^{b}}{\mathtt{a}_{t}^{l}r_{t}U_{t}-\mathtt{b}_{t}^{l}}&\text{if }r_{\text{min}}\leq r_{t},\text{ and }r_{t}U_{t}\geq r_{\text{max}},\\ 0&\text{if }r_{t}U_{t}>r_{\text{max}},\\ \end{cases}$
The objective of our protocol is to set the interest rate $r_{t}$ such that the utilization remains close to a desired utilization $U^{*}$. The optimal interest rate $r^{*}_{\text{utl}}$ that achieves this desired utilization can be derived as follows:  [^2] We develop similar results for the objective of maximizing revenue instead of optimizing rate deviation. See Appendix [A.2]
(15) $\frac{-\mathtt{a}_{t}^{b}\,r^{*}_{\text{utl}}+\mathtt{b}_{t}^{b}}{\mathtt{a}_{t}^{l}r^{*}_{\text{utl}}\,U^{*}-\mathtt{b}_{t}^{l}}=U^{*}\implies r^{*}_{\text{utl}}=\frac{\mathtt{b}_{t}^{b}+\mathtt{b}_{t}^{l}U^{*}}{\mathtt{a}_{t}^{b}+\mathtt{a}_{t}^{l}(U^{*})^{2}}$
@@
Initialize: $U^{*}$, $\zeta$ (randomness probability), $\xi$ (error threshold to do randomization) for each time step $t$ do
Read $\hat{\mathtt{a}}_{t}^{b},\hat{\mathtt{b}}_{t}^{b},\hat{\mathtt{a}}_{t}^{l},\hat{\mathtt{b}}_{t}^{l},\mathbf{P}^{l}_{t},\mathbf{P}^{b}_{t}$ from Algorithm 1
$\text{Var}(\hat{\mathtt{a}}_{t}^{b})\leftarrow\mathbf{P}^{b}_{t}(1,1)$, $\text{Var}(\hat{\mathtt{b}}_{t}^{b})\leftarrow\mathbf{P}^{b}_{t}(2,2)$, $\text{Var}(\hat{\mathtt{a}}_{t}^{l})\leftarrow\mathbf{P}^{l}_{t}(1,1)$, $\text{Var}(\hat{\mathtt{b}}_{t}^{l})\leftarrow\mathbf{P}^{l}_{t}(2,2)$
$\mathbb{E}\left[r_{t}\right]\leftarrow\frac{\hat{\mathtt{b}}_{t}^{b}+\hat{\mathtt{b}}_{t}^{l}U^{*}}{\hat{\mathtt{a}}_{t}^{b}+\hat{\mathtt{a}}_{t}^{l}(U^{*})^{2}}$
$\text{Var}(r_{t})\leftarrow\frac{1}{\big{(}\hat{\mathtt{a}}_{t}^{b}+\hat{\mathtt{a}}_{t}^{l}(U^{*})^{2}\big{)}^{2}}\left(\text{Var}(\hat{\mathtt{b}}_{t}^{b})+(U^{*})^{2}\text{Var}(\hat{\mathtt{b}}_{t}^{l})\right)$
$\quad+\frac{\big{(}\hat{\mathtt{b}}_{t}^{b}+\hat{\mathtt{b}}_{t}^{l}U^{*}\big{)}^{2}}{\big{(}\hat{\mathtt{a}}_{t}^{b}+\hat{\mathtt{a}}_{t}^{l}(U^{*})^{2}\big{)}^{4}}\left(\text{Var}(\hat{\mathtt{a}}_{t}^{b})+(U^{*})^{4}\text{Var}(\hat{\mathtt{a}}_{t}^{l})\right)$
      Sample $r_{t}\sim\mathbb{N}(\mathbb{E}\left[r_{t}\right],\text{Var}[r_{t}])$
end for
@@
with probability at least $1-\delta$, where $\limsup_{t\to\infty}\psi_{t}(\rho)=\frac{\nu^{2}(1-\rho^{N})^{2}}{\rho^{2N}\ln\left(\frac{1}{\rho}\right)}$ and $N$ is given by: $N=\Theta\left(\frac{\ln 1/\delta}{\ln\min_{t}\left(\text{diam}(\mathbf{P}^{b}_{t}+\mathbf{P}^{l}_{t})\right)}\right).$
@@
Adapting the forgetting factor The forgetting factor plays a crucial role in determining how quickly the RLS estimator reacts to changes in the underlying parameters. An optimally tuned $\rho(t)$ ensures higher stability and precision when parameters change slowly, while providing traceability when they change rapidly. Paleologu et al. (Paleologu et al., 2008) propose a method to adapt the forgetting factor based on the variance of the a posterior error, defined as $\varepsilon_{t}\coloneqq(y_{t}-\mathbf{x}_{t}^{T}\hat{\bm{\theta}}_{t})^{2}$. They argue that for a perfectly tuned filter, the error variance is determined solely by the observation noise i.e., $\text{Var}(\varepsilon_{t})=\hat{\sigma}(t),$ where $\hat{\sigma}(t)$ is the estimated noise variance. Using this relationship and the RLS update rules, the following expression for the adaptive forgetting factor is derived: $\rho(t)=\min\left\{\frac{\hat{\sigma}_{q}(t)\hat{\sigma}(t)}{\xi+\left|\hat{\sigma}_{e}(t)-\hat{\sigma}(t)\right|},\lambda_{\max}\right\},$ where $q_{t}\coloneqq\mathbf{x}_{t}^{T}\mathbf{P}_{t-1}\mathbf{x}_{t}$ and $e_{t}\coloneqq(y_{t}-\mathbf{x}_{t}^{T}\hat{\bm{\theta}}_{t-1})^{2}$ (the a priori error), and $\hat{\sigma}_{q}(t)$, $\hat{\sigma}_{e}(t)$ are their respective empirical variances. To estimate $\hat{\sigma}^{2}(t)$, the noise variance, a longer period of the a priori error signal $e(t)$ is used through an exponentially weighted moving average: $\hat{\sigma}^{2}(t)=\beta\hat{\sigma}^{2}(t-1)+(1-\beta)e^{2}(t).$
4.2. Risk parameters
@@
Intuitively, the lemma presented above provides a method for adjusting the liquidation threshold $LT_{t}$ and the collateral factor $c_{t}$ to minimize these risks. By setting $LT_{t}$ appropriately, the protocol can keep expected defaults below a threshold, while also ensuring that liquidations remain below a certain threshold. The relationship between the loan-to-value ratio and the liquidation threshold is key for dynamically adjusting risk parameters. This approach is based on (Bastankhah et al., 2024), where $LT_{t}$ serves as an upper bound for the loan-to-value ratio, helping to limit the likelihood of default.
5. Adversarial analysis
@@
5.1. Learning-based interest rate controller
Learning-based interest rate controllers estimate demand and supply while incorporating prior beliefs to set optimal rates. As adaptivity increases, these controllers become more vulnerable to manipulation, quickly updating beliefs based on a few manipulated samples. To quantify this impact, we analyze an adaptive learning-based algorithm, $\mathtt{LC}$, which alternates between exploration and exploitation phases. During exploration, the controller detects discrepancies between observed and expected demand and supply, testing various rates to update its beliefs. Once stabilized, the controller enters exploitation, setting rates based on the learned functions. This cycle repeats as new changes are detected. We model $\mathtt{LC}$ as an auction-like mechanism. In exploration, borrowers and lenders report their minimum (or maximum) acceptable rates and quantities. $\mathtt{LC}$ infers the demand and supply curves from this data. In exploitation, it sets a single rate $r_{t}$, allowing lenders with bids $r_{t}$ or lower to deposit, and borrowers with bids $r_{t}$ or higher to receive loans. This mimics a uniform price auction, where participants’ reported bids influence the interest rate, maximizing utility based on private valuations. Participant utility depends on whether the rate falls within their acceptable range and the portion of their supply or demand allocated to the protocol or external markets. More details on this auction implementation are provided in Appendix [A.3].